Let O be the set of odd numbers and O’ = {1, 5, 9, 13, 17, ...} be its subset. Define

the bijections, f and g as:

f : O "\to" O’, f(d) = 2d - 1, "\forall" d "\in" O.

g : "\Nu" "\to" O, g(n) = 2n + 1, "\forall" n "\isin" "\Nu" .

Using only the concept of function composition, can there be a bijective map from "\Nu"

to O’? If so, compute it. If not, explain in details why not.



1
Expert's answer
2022-02-22T15:10:50-0500

Given that the functions "f \text{~and~} g" are bijections, defined as "f(d) = 2d - 1, \forall d \in O," "g(n) = 2n + 1, \forall n \in \N". Here "O' = \{4x + 1 : x\in \N \cup \{0\}\} ."

But we see that, "g : \N \to O" is not surjective since there is no preimage of 1 in "\N".

For, "1= g(n)=2n+1 \implies n = 0", but "0\notin \N".

Hence there cannot be a bijective function from "\N \to O'".


For "g" must be bijective, it is must be defined as "g: \N \cup \{0\} \to O".

Setting "g: \N \cup \{0\} \to O" and using the fact "composition of two bijective functions is bijective", we have "h = f \circ g : \N \cup\{0\}\to O'" is also a bijective function. The composition of functions is

"h(x) = (f \circ g)(x) = f(g(x)) = f(2x+1) = 2(2x+1)-1 = 4x+1,~\forall x \in \N \cup \{0\}."